The Shirt Works sells a large variety of tee shirts and sweatshirts. Steve Hooper, the owner, is thinking of expanding his sales by hiring high school students, on a commission basis, to sell sweatshirts bearing the name and mascot of the local high school.

       These sweatshirts would have to be ordered from the manufacturer six weeks in advance, and they could not be returned because of the unique printing required. The sweatshirts would cost Hooper $21.00 each with a minimum order of 280 sweatshirts. Any additional sweatshirts would have to be ordered in increments of 50.

       Since Hooper’s plan would not require any additional facilities, the only costs associated with the project would be the costs of the sweatshirts and the costs of the sales commissions. The selling price of the sweatshirts would be $42.00 each. Hooper would pay the students a commission of $5.00 for each shirt sold.


Required:
1.

To make the project worthwhile, Hooper would require a $13,600 profit for the first three months of the venture. What level of unit sales and dollar sales would be required to reach this target net operating income? (Round your intermediate calculations to 2 decimal places.)

    

2.

Assume that the venture is undertaken and an order is placed for 280 sweatshirts. What would be Hooper’s break-even point in unit sales and in dollar sales? (Round your intermediate calculations to the nearest whole number. Round up your final break even answers to the nearest whole number.)
Answer #1
1
Sales level in units 850 sweatshirts =13600/(42-5-21)
Sales level in dollars 35700 =850*42
2
Break even point in units 159 sweatshirts =(280*21)/(42-5)
Break even point in sales dollars 6678 =159*42